At a coffee shop, 80% of customers purchase a drink, 40% purchase something to eat, and 20% purchase both a drink and something to eat. 3. (a) What fraction of customers purchase only a drink? (b) Given that someone bought something to eat, what is the probaility they also got a drink? (c) Given that someone bought a drink, what is the probability they didnt buy anything to eat?
Answer #1

Let D denote drink, E denote something to eat.

Given that P(D) = 0.80, P(E) = 0.40, P(D and E) = 0.20.

a. P(D and E') = P(D) - P(D and E) = 0.80-0.20 = 0.60

b. P(D|E) = P(D and E) / P(E) = 0.20/0.40 = 0.5

c. P(E'|D) = P(D and E') /P(D) = 0.60/0.80 = 0.75